the local coffee shop combines two kinds of beans mixture of beans. if the Ethiopian harrar beans cost $7.25 per pound, and Arabian mocha beans cost $13.70 per pound, how many pounds of each should be used to make 3 pounds of a blend that costs $10.26 per pound?
Answer by checkley79(3341)

7.25x+13.70(3-x)=10.26*3
7.25x+41.1-13.70x=30.78
6.45x=30.78-41.1
6.45x=10.32
x=10.32/6.45
x=1.6 pounds of Ethiopian is used.
3-1.6=1.4 pounds of Arabian is used.
Proof:
7.25*1.6+13.70*1.4=10.26*3
11.6+19.18=30.78
30/78=30.78